OpenStudy (anonymous):

Which of the following functions has a period of 4 π ? y = sin 1/4x y = sin 1/2x y = sin2x y = sin4x

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

the period of \[\sin(bx)\] is \[\frac{2\pi}{b}\]

OpenStudy (misty1212):

so not the first one there \(b=\frac{1}{4}\) and \[\frac{2\pi}{\frac{1}{4}}=2\pi\times 4=8\pi\] try the second one

OpenStudy (anonymous):

i got 4n so this would be my anwser

OpenStudy (anonymous):

im i right